# DreamAnywhere: Object-Centric Panoramic 3D Scene Generation

### v1

@@
Abstract.
@@
3D scene generation, text-to-scene, diffusion models
journal: TOGccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Graphics systems and interfaces
1. Introduction
@@
Existing text-to-3D approaches leverage the richness and stylistic diversity of internet-scale diffusion models for progressive outpainting (Höllein et al., 2023) or distillation sampling (Shriram et al., 2024; Li et al., 2025), but struggle to capture global coherence, leading to increasing semantic drift and structural collapse due to the difficulty of integrating depth estimates into a single 3D scene representation. Video diffusion models have been combined with large reconstruction models (Yu et al., 2024c; Chen et al., 2025b) but still face limitations in generating complete scenes observable from vastly different viewpoints maintaining coherency. Panoramic 360° images are an inherently good starting point for immersive viewing, in fact (Li et al., 2024; Zhou et al., 2024a) combine panoramic images with inpainting to enable parallax effects when moving away from the origin, further extended to multiple layers by (Yang et al., 2024c). Similarly, approaches designed for 2D fly-throughs (Liu et al., 2021; Fridman et al., 2024; Yu et al., 2024b, a) can create visually convincing animations, but lack the global 3D structure required for full scene exploration.
Whilst large-scale and artistic 3D scene data remains scarce, there exists a wealth of 3D data for individual objects, offering a valuable resource for compositional approaches. Moreover, decomposing a 3D environment into individual entities offers important advantages for downstream applications (Tulsiani et al., 2018; Armeni et al., 2019). If the layout and description of objects in a scene are known or provided manually, immersive scenes can be generated (Li et al., 2025). However, optimizing scene layouts end-to-end continues to be a major challenge (Zhang et al., 2023c; Epstein et al., 2024).
Building on these insights, we present DreamAnywhere, a text-to-3D generation framework that leverages 360° panoramic image generation to drive scene style and geometry, while decomposing scenes into individual objects to preserve structural coherence and enable the generation of navigable 3D environments.
Our main contribution is a unified, modular system for 3D scene generation, enabled by the following key components:
  • A panoramic image generation method that jointly fine-tunes a 360° image diffusion model with a perspective conditioning mechanism, improving out-of-domain sampling.
@@
Figure 2. An overview of our architecture. We start by generating a 360° panorama, followed by instance segmentation to separate the foreground objects from the scene background. Segmented objects are lifted to 3D using semantic and geometric priors, while a 3D inpainting process converts the image to a 3DGS representation and fills disoccluded regions. Finally, the scene is composited, enabling navigation of the 3D environment at interactive rates.
2. Related Work
@@
Our 3D scene generation work lies at the intersection of image-based generation and compositional scene synthesis, bridging recent advances in both areas. For an in-depth discussion we refer readers to recent surveys on scene generation  (Li et al., 2023; Tang et al., 2025) and object generation  (Li et al., 2023; Po et al., 2024), and below we highlight the key developments most relevant to our approach.
2.1. Image-based Scene Generation
@@
A successful line of work in image-based scene generation focuses on monolithic representations, where the entire scene is captured in a single unified 3D structure. A common technique involves iterative outpainting with image diffusion models to progressively extend the scene along a camera trajectory (Fridman et al., 2024; Höllein et al., 2023; Yu et al., 2024b, a; Liang et al., 2024). Although effective to some extent, these methods often suffer from the accumulation of visual artifacts and distortions across iterations, and offer limited flexibility for isolating and manipulating individual scene components.
@@
An alternative set of techniques reconstructs 3D scenes from 2D images using inpainting-based methods, which lift partial views into 3D and fill occluded regions through inpainting (Shriram et al., 2024; Zhang et al., 2023b). Although these reconstructions can appear photorealistic from specific angles, their quality tends to degrade when viewed from arbitrary viewpoints.
@@
More recently, panorama-based approaches such as PanoDreamer (Paliwal et al., 2024) and PanoDiffusion (Wu et al., 2023b) attempt to synthesize entire 360° panoramic images with depth information through in- or outpainting. Rather than iteratively expanding standard 2D images, other methods focus on creating 360° panorama images in a single step (Tang et al., 2024; Li et al., 2024; Zhou et al., 2024a; Stan et al., 2023). Another area of research addresses the handling of occlusions via multi-view generation (Ye et al., 2024), layered approaches (Yang et al., 2024c) or explicitly defining the corresponding entities (Zhou et al., 2025). Video diffusion models have strong temporal coherency but are still limited in generating 3D content that is consistent along long camera trajectories  (Yu et al., 2024c; Ren et al., 2025; Chen et al., 2025b; Zhai et al., 2025; Zhang et al., 2025).  While image-based methods offer strong visual quality and global coherence, they generally lack fine-grained control over individual scene elements and are limited in navigational flexibility, especially when compared to compositional scene generation approaches, which we discuss in the following section.
2.2. Compositional Scene Generation
@@
In contrast to monolithic approaches, compositional scene generation focuses on constructing scenes from multiple, independently generated objects which can be edited or animated. Recent advances in this area predominantly leverage pre-trained image diffusion models for scene generation (Huang et al., 2024; Dogaru et al., 2024; Zhou et al., 2024b; Han et al., 2024; Chen et al., 2025a). The key differences between these methods lie in the strategies used to layout the scene.
@@
Image-guided methods (Huang et al., 2024; Dogaru et al., 2024; Zhang et al., 2023c; Zhou et al., 2024b; Han et al., 2024) use generated or real images to infer each object’s location and orientation in the scene. Similarly, distillation-based approaches (Epstein et al., 2024; Li et al., 2025; Chen et al., 2025a) employ diffusion models to guide object placement. However, since these methods typically rely on single (non-panoramic) images, they are often limited in terms of scale and global coherence of the generated scenes.
@@
To overcome these limitations, proxy-guided techniques introduce explicit control over scene structure, for example by letting users define the scene structure through various semantic annotations or by placing proxy geometries (Wu et al., 2024; Schult et al., 2024; Fang et al., 2023; Cohen-Bar et al., 2023; Yang et al., 2024b). While these techniques provide great flexibility in terms of scale and coherence, they typically require extensive user input and manual scene design.
@@ 3. Method
The DreamAnywhere architecture proceeds through three stages. In the first stage, a $360\degree{}$ panoramic image $I_{F}$ is generated from a text prompt and stylized by an input perspective image. From $I_{F}$, foreground objects are identified, segmented, and erased, producing a clean background image $I_{B}$. Then, the depth of $I_{B}$ is aligned to that of $I_{F}$, producing a single, unified depth map and laying the groundwork for subsequent 3D composition. The second stage focuses on generating high-fidelity 3D Gaussian Splatting (3DGS) models of the segmented objects and computing their transformations for accurate placement in the 3D scene. In parallel, the third stage converts the background image $I_{B}$ into a high-resolution 3DGS point cloud. During this process, disoccluded regions are inpainted to ensure visual consistency across viewpoints under novel camera offsets. Finally, the outputs of the second (objects) and third (background) stages are fused into a unified 3DGS scene, enabling immersive and interactive navigation. An overview of our pipeline is illustrated in Figure 2. The remainder of this section presents all the components of our system, emphasizing the most essential technical and implementation details. Additional information and extended ablation studies regarding 2D and 3D inpainting, object and background generation, pose estimation and scene compositing are available in the supplementary material.
3.1. Image Generation
3.1.1. 360° Panorama Diffusion
@@ 3.1. Image Generation
The initial stage of our pipeline generates the 360° panorama image $I_{F}$ from a user-provided text prompt. To improve generalization and style fidelity, we guide the generation process using a perspective image derived from the same prompt, providing soft conditioning without enforcing pixel-level alignment. Existing panoramic diffusion models are typically LoRA-fine-tuned on small, less-diverse equirectangular datasets, limiting their ability to generalize to out-of-distribution styles/compositions despite these capabilities existing in the pre-trained base model.
To address this, we condition the diffusion process of the 360° image with a perspective image derived from the same text prompt, using an IP-Adapter-style mechanism (Ye et al., 2023). We observe that simply adding this mechanism post-hoc to a panoramic LoRA model is ineffective due to a mismatch in perspective (for IP-Adapter) and panoramic (for LoRA) training distributions as shown in the ablation in Figure 9. Instead, we jointly fine-tune the panoramic LoRA with the IP-adapter, enabling effective style transfer from perspective to panoramic images. We base our model on Stable-Diffusion-3.5-Large and fine-tune a rank-8 LoRA on top of all attention blocks on a mix of indoor (Chang et al., 2017) and outdoor (PolyHaven, 2025). We train the model for 10k iterations with a learning rate of $10^{-5}$ and a batch size of 4.
3.1.2. Instance Segmentation
@@
To identify suitable candidates for object reconstruction in the $360\degree{}$ images, we employ the off-the-shelf open-set segmentation pipeline Grounded-SAM (Ren et al., 2024). To improve robustness in stylized or out-of-distribution scenes, we replace the original class identification stage (RAM (Zhang et al., 2023a)) with a larger vision-language model (GPT-4V), improving category recognition under diverse visual conditions. Given the limitations of current object generators, we exclude categories that are not good candidates for object reconstruction and are typically considered background elements (such as grass, walls, and floors). To further ensure high-quality segmentation, we sort all detected objects by their predicted IoU confidence and retain only the highest-scoring masks. For each retained object, the pipeline outputs a class identifier, a textual description, and a fine-grained segmentation mask, which are later used in Section 3.2.
@@
Figure 3. Our image generation pipeline uses a diffusion model to denoise an equirectangular image (Yang et al., 2024c; Feng et al., 2023) but additionally combines perspective image conditioning through a decoupled cross-attention layer (Ye et al., 2023). We train an equirectangular LoRA jointly with the perspective pre-trained cross attention layer, by encoding random perspective crops of the panorama, showing stronger generalization capabilities to out-of-domain panorama sampling.
3.1.3. 2D Inpainting
@@
We generate the background panorama $I_{B}$ by erasing the foreground objects identified in the full panorama using the union of their segmentation masks. To synthesize plausible new content in the masked regions, we employ a diffusion-based inpainting model, Dreamshaper-8-Inpainting,  (Lykon, 2023), chosen for its ability to synthesize meaningful background content. We reduce hallucinations and artifacts in a two-stage process. First, we pre-inpaint the masked regions using LaMA (Suvorov et al., 2022) to obtain a coherent object-free initialization. Second we feed this through a custom LoRA fine-tuned on panoramic images from the Structured3D dataset(Zheng et al., 2020) for indoor defurnishing (in a similar spirit to (Slavcheva et al., 2024)). To further improve image quality, we apply standard techniques: the segmentation masks are slightly dilated, wrap-around consistency is enforced via horizontal padding (256 pixels), and Poisson blending seamlessly integrates the original panorama with the newly inpainted content. Although we experimented with repurposing the panoramic LoRA for inpainting, we found that a dedicated, specialized pipeline significantly reduces hallucinations without requiring manual intervention, especially in indoor scenes. The resulting empty panorama image is then lifted to 3D to form the background point cloud as described in Section 3.3.
3.1.4. Depth Estimation
@@
We employ different depth estimation strategies for indoor and outdoor scenes due to their distinct geometric challenges. For indoor scenes, we use EGformer (Yun et al., 2023), which excels in predicting the axis-aligned structure typical of indoor environments. For outdoor scenes, we instead use the 360MonoDepth framework (Rey-Area et al., 2022), which aligns multiple perspective depth estimates obtained with Depth-Anything-V2 and possesses strong zero-shot generalization capabilities. To address unreliable depth predictions in low-detail regions, such as the sky, we explicitly detect and segment these areas using GroundedSAM (Ren et al., 2024) and assign the maximum depth value to ensure accurate representation. By construction, we obtain two independent depth estimates for $I_{F}$ and $I_{B}$ and perform an alignment step from $I_{B}$ to $I_{F}$ to ensure that objects are placed properly in their corresponding locations. Similarly to concurrent work, we find that a global affine transform is not sufficient and instead employ a per-pixel shift and global scale that minimizes the $L2$ depth error in overlapping regions between the images, subject to as-rigid-as-possible constraints (Sorkine and Alexa, 2007).
3.2. High Quality Object Generation
@@
To enable stylistically aligned and geometrically consistent 3D object reconstructions, we condition image-to-3D models using RGBD data derived from the segmented panorama. However, segmented object images are often incomplete, low-resolution, visible from suboptimal viewpoints, or contain generative artifacts (such as distortions or incorrect limbs). To overcome these problems, we introduce a resynthesis step that produces a high-quality reference image of the object, utilizing both semantic and geometric cues. This image is then used to guide the 3D reconstruction of each object. Our pipeline consists of three stages: reference image generation, 3D object generation, and pose estimation for accurate placement in the scene. We illustrate this process in Figure 4.
@@
Figure 4. Our object reconstruction pipeline leverages the panorama and style information to generate a high-resolution reference image to be used for multi-view generation. The generated multi-view images are then transformed into 3D Gaussian splats through a reconstruction pipeline. Finally, we align the generated object with the original and place it in the scene.
3.2.1. Reference Image Generation

### v2

@@
DreamAnywhere: Object-Centric Panoramic 3D Scene Generation
Edoardo A. Dominici1  Jozef Hladký1  Floor Verhoeven1  Lukas Radl2
Thomas Deixelberger12  Stefan Ainetter1  Philipp Drescher12  Stefan Hauswiesner1
Arno Coomans1  Giacomo Nazzaro1  Konstantinos Vardis1  Markus Steinberger1 2
1Huawei Technologies   2Graz University of Technology
Figure 1: Starting from a text prompt, our method generates 3D scenes leveraging a 360° panoramic image as intermediate representation to extract, reconstruct and compose objects in the environment. Our scenes, represented as 3D Gaussian splats, support long-range exploration and have high structural coherence even under large camera offsets. This improves immersive scene generation for existing applications, enables intuitive object-level editing and makes them strong 3D priors for world-to-world transfer models [1].
@@
1 Introduction
@@
Existing text-to-3D approaches leverage the richness and stylistic diversity of internet-scale diffusion models for progressive outpainting [18] or distillation sampling [40, 25], but struggle to capture global coherence, leading to increasing semantic drift and structural collapse due to the difficulty of integrating depth estimates into a single 3D scene representation. Video diffusion models have been combined with large reconstruction models [60, 5] but still face limitations in generating complete scenes observable from vastly different viewpoints maintaining coherency. Panoramic 360° images are an inherently good starting point for immersive viewing, in fact [8, 69] combine panoramic images with inpainting to enable parallax effects when moving away from the origin, further extended to multiple layers by [53]. Similarly, approaches designed for 2D fly-throughs [27, 14, 59, 58] can create visually convincing animations, but lack the global 3D structure required for full scene exploration.
Whilst large-scale and artistic 3D scene data remains scarce, there exists a wealth of 3D data for individual objects, offering a valuable resource for compositional approaches. Moreover, decomposing a 3D environment into individual entities offers important advantages for downstream applications [47, 3]. If the layout and description of objects in a scene are known or provided manually, immersive scenes can be generated [25]. However, optimizing scene layouts end-to-end continues to be a major challenge [65, 11].
Building on these insights, we present DreamAnywhere, a text-to-3D generation framework that leverages 360° panoramic image generation to drive scene style and geometry, while decomposing scenes into individual objects to preserve structural coherence and enable the generation of navigable 3D environments. We make the following contributions:
  • A novel unified and modular system for generating 3D scenes with high 3D coherence and visual fidelity while facilitating interactivity.
  • A panoramic image generation method that jointly fine-tunes a 360° image diffusion model with a perspective conditioning mechanism, improving out-of-domain sampling for creative applications.
@@
Figure 2: An overview of our architecture. We start by generating a 360° panorama, followed by instance segmentation to separate the foreground objects from the scene background. Segmented objects are lifted to 3D using semantic and geometric priors, while a 3D inpainting process converts the image to a 3DGS representation and fills disoccluded regions. Finally, the scene is composited, enabling navigation of the 3D environment at interactive rates.
2 Related Work
@@
Our 3D scene generation work lies at the intersection of image-based generation and compositional scene synthesis, bridging recent advances in both areas. For an in-depth discussion we refer readers to recent surveys on scene generation  [24, 46] and object generation  [24, 32].
2.1 Image-based Scene Generation
@@
A successful line of work in image-based scene generation focuses on monolithic representations, where the entire scene is captured in a single unified 3D structure. A common technique involves iterative outpainting with image diffusion models to progressively extend the scene along a camera trajectory [14, 18, 59, 58, 26]. Although effective to some extent, these methods often suffer from the accumulation of visual artifacts and distortions across iterations, and offer limited flexibility for isolating and manipulating individual scene components.
@@
An alternative set of techniques reconstructs 3D scenes from 2D images using inpainting-based methods, which lift partial views into 3D and fill occluded regions through inpainting [40, 64]. Although these reconstructions can appear photorealistic from specific angles, their quality tends to degrade when viewed from arbitrary viewpoints.
@@
More recently, panorama-based approaches such as PanoDreamer [31] and PanoDiffusion [50] attempt to synthesize entire 360° panoramic images with depth information through in- or outpainting. Rather than iteratively expanding standard 2D images, other methods focus on creating 360° panorama images in a single step [45, 8, 69, 43]. Another area of research addresses the handling of occlusions via multi-view generation [57], layered approaches [53] or explicitly defining the corresponding entities [71]. Video diffusion models have strong temporal coherency but are still limited in generating 3D content that is consistent along long camera trajectories  [60, 36, 5, 62, 66, 19]. While image-based methods offer strong visual quality and global coherence, they generally lack fine-grained control over individual scene elements and are limited in navigational flexibility, especially when compared to compositional scene generation approaches, which we discuss in the following section.
2.2 Compositional Scene Generation
@@
In contrast to monolithic approaches, compositional scene generation focuses on constructing scenes from multiple, independently generated objects which can be edited or animated. Recent advances in this area predominantly leverage pre-trained image diffusion models for scene generation [20, 2, 70, 16, 6]. The key differences between these methods lie in the strategies used to layout the scene.
@@
Image-guided methods [20, 2, 65, 70, 16] use generated or real images to infer each object’s location and orientation in the scene. Similarly, distillation-based approaches [11, 25, 6] employ diffusion models to guide object placement. However, since these methods typically rely on single (non-panoramic) images, they are often limited in terms of scale and global coherence of the generated scenes.
@@
To overcome these limitations, proxy-guided techniques introduce explicit control over scene structure, for example by letting users define the scene structure through various semantic annotations or by placing proxy geometries [51, 38, 12, 9, 54]. While these techniques provide great flexibility in terms of scale and coherence, they typically require extensive user input and manual scene design.
@@ 3 Method
The DreamAnywhere architecture proceeds through three stages as illustrated in Figure 2. In the first stage, a $360\degree{}$ panoramic image $I_{F}$ is generated from a text prompt and stylized by an input perspective image. From $I_{F}$, foreground objects are identified, segmented, and erased, producing a clean background image $I_{B}$. Then, the depth of $I_{B}$ is aligned to that of $I_{F}$, producing a single, unified depth map and laying the groundwork for subsequent 3D composition. The second stage focuses on generating high-fidelity 3D Gaussian Splatting (3DGS) models of the segmented objects and computing their transformations for accurate placement in the 3D scene. In parallel, the third stage converts the background image $I_{B}$ into a high-resolution 3DGS point cloud. During this process, disoccluded regions are inpainted to ensure visual consistency across viewpoints under novel camera offsets. Finally, the outputs of the second (objects) and third (background) stages are fused into a unified 3DGS scene, enabling immersive and interactive navigation. The remainder of this section presents all the components of our system, emphasizing the most essential technical and implementation details. Additional information and extended ablation studies regarding 2D and 3D inpainting, object and background generation, pose estimation and scene compositing are available in the supplementary material.
3.1 Image Generation
3.1.1 360° Panorama Diffusion
@@ 3.1 Image Generation
The initial stage of our pipeline generates the 360° panorama image $I_{F}$ from a user-provided text prompt. Existing text to panorama diffusion models are typically LoRA-fine-tuned on small, less-diverse equirectangular datasets, limiting their ability to generalize to out-of-distribution styles/compositions despite these capabilities existing in the pre-trained base model. To improve generalization and style fidelity, we guide the generation process using a perspective image derived from the same prompt, providing soft conditioning without enforcing pixel-level alignment. We achieve this by using an IP-Adapter-style mechanism [56] that introduces separate cross-attention layers in all transformer blocks of the diffusion model. We observe that simply adding this mechanism post-hoc to a panoramic LoRA model is ineffective due to a mismatch in perspective (for IP-Adapter) and panoramic (for LoRA) training distributions as shown in the ablation in Figure 8. Instead, we jointly fine-tune the panoramic LoRA with the IP-adapter using random perspective renders of the equirectangular image (Figure 3), enabling effective style transfer from perspective to panoramic images.
We base our model on Stable-Diffusion-3.5-Large and fine-tune a rank-8 LoRA on top of all attention blocks on a mix of indoor [4] and outdoor [33] images. We train the model for 10k iterations with a learning rate of $10^{-5}$ and a batch size of 4.
3.1.2 Instance Segmentation
@@
To identify suitable candidates for object reconstruction in the $360\degree{}$ images, we employ the off-the-shelf open-set segmentation pipeline Grounded-SAM [35]. To improve robustness in stylized or out-of-distribution scenes, we replace the original class identification stage (RAM [67]) with a larger vision-language model (GPT-4V), improving category recognition under diverse visual conditions. Given the limitations of current object generators, we exclude categories that are not good candidates for object reconstruction and are typically considered background elements (such as grass, walls, and floors). To further ensure high-quality segmentation, we sort all detected objects by their predicted IoU confidence and retain only the highest-scoring masks. For each retained object, the pipeline outputs a class identifier, a textual description, and a fine-grained segmentation mask, which are later used in Section 3.2.
@@
Figure 3: Our image generation pipeline uses a diffusion model to denoise an equirectangular image [53, 13], but additionally combines perspective image conditioning through a decoupled cross-attention layer [56]. We train an equirectangular LoRA jointly with the perspective pre-trained cross attention layer, by encoding random perspective crops of the panorama, showing stronger generalization capabilities to out-of-domain panorama sampling.
3.1.3 2D Inpainting
@@
We generate the background panorama $I_{B}$ by erasing the foreground objects identified in the full panorama using the union of their segmentation masks. To synthesize plausible new content in the masked regions, we employ a diffusion-based inpainting model, Dreamshaper-8-Inpainting [29], chosen for its ability to synthesize meaningful background content. We reduce hallucinations and artifacts in a two-stage process. First, we pre-inpaint the masked regions using LaMA [44] to obtain a coherent object-free initialization. Second we feed this through a custom LoRA fine-tuned on panoramic images from the Structured3D dataset [68] for indoor defurnishing (in a similar spirit to [41]). To further improve image quality, we apply standard techniques: the segmentation masks are slightly dilated, wrap-around consistency is enforced via horizontal padding (256 pixels), and Poisson blending seamlessly integrates the original panorama with the newly inpainted content. Although we experimented with repurposing the panoramic LoRA for inpainting, we found that a dedicated, specialized pipeline significantly reduces hallucinations without requiring manual intervention, especially in indoor scenes. The resulting empty panorama image is then lifted to 3D to form the background point cloud as described in Section 3.3.
3.1.4 Depth Estimation
@@
We employ different depth estimation strategies for indoor and outdoor scenes due to their distinct geometric challenges. For indoor scenes, we use EGformer [61], which excels in predicting the axis-aligned structure typical of indoor environments. For outdoor scenes, we instead use the 360MonoDepth framework [37], which aligns multiple perspective depth estimates obtained with Depth-Anything-V2 and possesses strong zero-shot generalization capabilities. To address unreliable depth predictions in low-detail regions, such as the sky, we explicitly detect and segment these areas using GroundedSAM [35] and assign the maximum depth value. By construction, we obtain two independent depth estimates for $I_{F}$ and $I_{B}$ and perform an alignment step from $I_{B}$ to $I_{F}$ to ensure that objects are placed properly in their corresponding locations. Similarly to concurrent work, we find that a global affine transform is not sufficient and instead employ a per-pixel shift and global scale that minimizes the $L2$ depth error in overlapping regions between the images, subject to as-rigid-as-possible constraints [42].
3.2 High Quality Object Generation
@@
To enable stylistically aligned and geometrically consistent 3D object reconstructions, we condition image-to-3D models using RGBD data derived from the segmented panorama. However, segmented object images are often incomplete, low-resolution, seen from suboptimal viewpoints, or contain generative artifacts (such as distortions or incorrect limbs). To overcome these problems, we introduce a resynthesis step that produces a high-quality reference image of the object, utilizing both semantic and geometric cues. This image is then used to guide the 3D reconstruction of each object. Our pipeline consists of three stages: reference image generation, 3D object generation, and pose estimation for accurate placement in the scene. We illustrate this process in Figure 4.
@@
Figure 4: Our object reconstruction pipeline leverages the panorama and style information to generate a high-resolution reference image to be used for multi-view generation. The generated multi-view images are then transformed into 3D Gaussian splats through a reconstruction pipeline. Finally, we align the generated object with the original and place it in the scene.
3.2.1 Reference Image Generation